[figure: a^{L}; a; a^{L}; b; a; b^{R}; b; b^{R}] \begin{aligned}{[a, b]^{1} \atop a} & \begin{aligned}{[a, b]^{1} \atop b} & & b↦c \\ & c \end{aligned} \\ & a, c \end{aligned} ; a, b↦a, c ; 1

tmp July 12, 2008 15:55